# PhreshPhish: A Real-World, High-Quality, Large-Scale Phishing Website Dataset and Benchmark

### v1

@@ 4 PhreshPhish Test and Benchmark Datasets
In contrast to our approach, we found most publicly available phishing website datasets either lack a test or benchmark dataset, or suffer from significant training-to-test set leakage, leading to inflated performance evaluation. Further, the base rate in most datasets is too high to serve as realistic benchmarks.
4.1 Test Dataset
Starting from the clean dataset, we first perform temporal partitioning, followed by leakage filtering to remove test data points that show high similarity to any point in the train set. The pseudocode to generate the test set is shown in Algorithm 2 in Appendix D.
@@ 4.1 Test Dataset
Leakage filter While temporal separation provides some protection against data leakage from the train to the test set, very similar points could still exist between train and test datasets. To mitigate this problem, we traverse the test set and discard points whose TF-IDF cosine distance is less than a specified threshold from any train points. LSH is used to make searching the train set more scalable. Figure 5 in Appendix D shows the temporal distribution of the train and test data sets based on date of collection.
4.2 PhreshPhish Benchmarks
@@
To encourage the development of more robust and generalizable phishing detection models, we provide a suite of benchmark datasets. The benchmark datasets are derived from the test dataset but are further processed to make them more challenging and representative of real-world scenarios. The steps to create the benchmarks are shown in Algorithm 3 in Appendix D.
Diversity enhancement We prune as well as augment the test dataset to make it more diverse and representative. We augment the benign data points with points corresponding to the top targeted brands. The targeted brands distribution is estimated from the phishing test data, and benign URLs are obtained using Google search. Similarly, one can augment the dataset with pages corresponding to top industry categories.
@@
Many data points have dissimilar URLs but very similar HTMLs. We prune such phishing data points from the test set. This is implemented similar to the leakage filter but with the test set and only the HTML is compared.
@@
Difficulty filter Applying the difficulty filter prunes out “easy” points, making the dataset more challenging to classify. There are several approaches to discovering easy points including exploring various data features, determining the most confounding feature values for the output classes, and selecting data points with those values. We take a simpler approach and assume the classification probability of a correctly classified data point can be taken as a proxy of its difficulty. We run the test set through a classifier trained on the training set. Correctly classified phishing points with score greater than $1-\delta$ are considered easy and candidates for pruning. We randomly sample and drop $k\%$ of the entire dataset. Specifically, we set $\delta=0.15$, and $k=10$. While we currently sample the candidates uniformly, an alternate approach would be to assign weights to the entire test set based on their scores, with higher weights for “easier” points, and then perform weighted sampling over the entire set. Figure 3 shows impact of applying the difficulty filter on performance.
@@
Base rate adjustment The base rate of phishing points in our test set is about 33%. Other publicly available test datasets have similar high base rates. However, in practice the base rates a detector observes are much lower, and vary depending on a multitude of factors including data provenance and the model pipeline used in deployment. To represent a wide range of scenarios, we create a large number of benchmark datasets with varying base rates. In particular, we create benchmark datasets with five different base rates varying from 0.05% to 5%. To account for variance, multiple instances, $m$, are sampled for each base rate, resulting in 404 total datasets. For lower base rates, the number of phishing points in a benchmark set are miniscule, and hence the variance is high.
@@
Determining $m$: How many instances, $m$, are required, given a base rate, $b$, corresponding to $k$ phishing points in a dataset of size $N$, for the model error estimate, $M_{e}$, to be accurate within a margin of error $e$ (e.g., ±$\pm 1\%$)? Since model error is an average, we can assume $M_{e}\sim\mathcal{N}(\mu_{M_{e}},\,\sigma_{M_{e}})$ (per central limit theorem). $m$ is sufficiently large when $m\geq\frac{z^{2}\cdot\hat{\sigma_{M_{e}}}^{2}}{e^{2}}$ where $\hat{\sigma_{M_{e}}}^{2}$ is the empirical variance of the model error and $z$ is the standard normal z-score corresponding to a desired confidence level. For each $b$, we sample multiple small sets of instances to compute average empirical variance, which is then used to estimate $m$. This analysis assumes sampling with replacement; sampling without replacement generally requires fewer samples, however, the difference is negligible when $k\ll N$, as is true in our case. Another advantage of sampling without replacement is that all available points are covered. In our experiments, we adopt sampling without replacement with $m$’s higher than the estimated lower bound with 1% margin of error and 95% confidence level. The base rates and the corresponding $m$’s in the benchmark datasets are: 0.05% (241), 0.1% (124), 0.5% (25), 1% (12), 5% (2).
@@
As shown in Figure 3, model performance is highly sensitive to the base rate, with performance improving as the base rate increases. Importantly, most of the strong performance results reported in the literature on this task can be largely attributed to the use of unrealistically high base rates.
Figure 3: Left: Precision-recall curves on the benchmark datasets. Right: Precision-recall curves on benchmark datasets with and without applying the difficulty filter
5 Baseline Models
@@
Phishing detection is a binary classification problem where the input space $\mathcal{X}$ is the Cartesian product $\mathcal{X}=\mathcal{U}\times\mathcal{H}$ where $\mathcal{U}$ denotes the space of URLs and $\mathcal{H}$ denotes the space of HTML content 333 Other input modalities such as host-based features (i.e. domain age, SSL certificate validity, etc.) could be used as well but we exclude them here since their inclusion tends to introduce latency that would be unacceptable for real-time detection.. The goal is to learn a function $f:\mathcal{U}\times\mathcal{H}\rightarrow[0,1]$ that estimates the conditional probability $P(Y=1|x),$ where $Y\in\{0,1\}$ is the binary class label (with $Y=1$ indicating a phishing page and $Y=0$ indicating a benign page) and $x=(u,h)\in\mathcal{X}$. The function is learned from a labeled dataset $\mathcal{D}=\{(x_{i},y_{i})\}^{N}_{i=1}$.
We implemented several representative approaches for real-time phishing detecting, including: (1) a linear model trained on n-gram and hand-crafted features; (2) a shallow feedforward neural network (FNN) using the same feature set; (3) a BERT-based model with a classifier head fine-tuned on raw HTML and URLs; and, (4) a large language model (LLM) used in a zero-shot prediction setting. For the linear and FNN models, input features include tokenized character-level n-grams from both the URL and HTML content, as well as hand-engineered features such as the length of the URL, number of nodes in the HTML DOM tree, and other domain-specific features. In contrast, the BERT and LLM-based approaches operate directly on raw textual inputs to learn rich representations. Our objective is not to optimize for state-of-the-art performance, but rather to establish baseline results on the benchmark datasets using common solution approaches. Model architecture and training details for each approach are described in Appendix E.
Linear model The linear model is a simple linear classifier $f(x)=W\cdot x+b$ implemented using LinearSVC from the scikit-learn library scikit-learn where parameters $W$ and $b$ are learned by minimizing the L2-regularized squared hinge loss: $\mathcal{L}(W,b)=\frac{1}{N}\sum_{i=1}^{N}(\max(0,1-y_{i}(W\cdot x_{i}+b)))^{2}+\frac{\lambda}{2}\|W\|^{2}_{2}$.
FNN model The FNN model uses the same input representation as the linear model but replaces the linear scoring function with a shallow neural network comprising of a single hidden layer with 32 neurons and ReLU activation. Input features are binarized and subject to frequency-based feature selection to reduce dimensionality. The model outputs a scalar probability estimate $f(x)=\sigma(W_{2}\cdot\texttt{ReLU}(W_{1}\cdot x+b_{1})+b_{2})$. This architecture introduces a small amount of non-linearity while maintaining tractable training and inference costs. Training is performed using binary cross-entropy loss with the Adam optimizer at a learning rate of 0.001.
BERT-based feature learning In this approach, we tokenize the raw, concatenated URL $u\in\mathcal{U}$ and HTML $h\in\mathcal{H}$ and input to a BERT-based encoder model, which generates a hidden contextualized representation, $\mathbf{h}=\texttt{Encoder}(\texttt{tokenize}(u\oplus h))\in\mathbb{R}^{n\times d}$, where $n$ is the number of tokens and $d$ is encoder model dimension. In particular, we use the gte-large pre-trained embedding model li2023towards , which computes the dense embedding as $x=\frac{1}{n}\sum_{i=1}^{n}h_{i}\in\mathbb{R}^{d}$. We attach a linear classification head for phishing detection, $f(x)=\sigma(W\cdot x+b)\in[0,1]$. We fully fine-tune the model using a binary cross-entropy loss for two epochs on the training dataset.
LLM-based prediction We prompt an LLM with a webpage URL and its procressed HTML to produce a rating indicating the likelihood that the page is a phishing site. The prompt is zero-shot, and the HTML is pruned to remove non-informative elements such as JavaScript, CSS and comments, inspired by koide2024chatphishdetector . This reduces input tokens and filters out content that is likely to be irrelevant to phishing detection. The returned rating is normalized and perturbed with small Gaussian noise to enable construction of a smooth precision-recall curve. Due to cost considerations and goal of establishing baseline performance, we use a smaller, earlier-generation model: gpt-4o-mini. We expect a larger, state-of-the-art reasoning LLM would produce improved performance. Further details, including the full prompt, are provided in Appendix E.
5.1 Results comparison and discussion
Users downstream of phishing detection systems are typically highly sensitive to false positives – instances where access to legitimate websites may be blocked. This is quantified by precision, the fraction of times users are correctly blocked. Consequently, commercial products aspire to a very high value of precision while also maintaining a good recall. Accordingly, we evaluate our models on two performance measures – the average precision, which is an approximation of the area under the precision-recall curve and the precision at a recall of 0.9.
We evaluate the models on the test dataset and the benchmark datasets for the five base rates. As expected, most models perform well on the test data set, which has a base rate of 18.3%. Table 2 summarizes the results for the four models. For each base rate, results are averaged over all corresponding benchmarks datasets. We find performance drops significantly as the base rate decreases. Overall, the GTE model performs the best. More evaluation results, including precision-recall curves and confusion matrices at multiple threshold values, are provided in Appendix E.
Model 0.05 0.10 0.50 1.0 5.0 18.3 (full)
Linear AP 0.5712 $\pm$ 0.0118 0.6592 $\pm$ 0.0134 0.8277 $\pm$ 0.0089 0.8745 $\pm$ 0.0052 0.9502 $\pm$ 0.0004 0.9940
P@R = 0.90 0.1232 $\pm$ 0.0173 0.2039 $\pm$ 0.0312 0.4179 $\pm$ 0.0528 0.5349 $\pm$ 0.0300 0.8684 $\pm$ 0.0111 0.9510
FNN AP 0.2395 $\pm$ 0.0097 0.3512 $\pm$ 0.0100 0.6415 $\pm$ 0.0104 0.7367 $\pm$ 0.0094 0.8787 $\pm$ 0.0025 0.9546
P@R = 0.90 0.0347 $\pm$ 0.0065 0.0625 $\pm$ 0.0115 0.1374 $\pm$ 0.0156 0.2176 $\pm$ 0.0357 0.6224 $\pm$ 0.0123 0.9035
GTE AP 0.7771 $\pm$ 0.0128 0.8142 $\pm$ 0.0101 0.8958 $\pm$ 0.0067 0.9236 $\pm$ 0.0065 0.9715 $\pm$ 0.0002 0.9918
P@R = 0.90 0.$2$$\pm$ 0.0296 0.3118 $\pm$ 0.0364 0.5666 $\pm$ 0.0353 0.7099 $\pm$ 0.0461 0.9322 $\pm$ 0.0034 0.9855
LLM AP 0.0223 $\pm$ 0.0012 0.0367 $\pm$ 0.0018 0.1306 $\pm$ 0.0061 0.2248 $\pm$ 0.0085 0.5552 $\pm$ 0.0139 0.7976
P@R = 0.90 0.0062 $\pm$ 0.0005 0.0123 $\pm$ 0.0012 0.0450 $\pm$ 0.0108 0.0690 $\pm$ 0.0138 0.2690 $\pm$ 0.0451 0.7780
@@ 5.1 Results comparison and discussion
Table 2: Mean average precision (AP) and mean precision at recall = 0.90 for each benchmark base rate, broken out by model. Values are reported as mean $\pm$ error, where the error reflects the maximum absolute deviation from the 95% bootstrap confidence interval.
6 Conclusions

### v2

@@ 4. Test and Benchmark Datasets
—David Patterson, 2017 Turing award winner (patterson2012better)
@@
In contrast to our approach, we found most publicly available phishing website datasets either lack a test or benchmark dataset, or suffer from significant training-to-test set leakage, leading to inflated performance evaluation. Further, the base rate in most datasets is too high to serve as realistic benchmarks. For example, Phishpedia (lin2021phishpedia), ChatPhishDetector (koide2024chatphishdetector), and PhishAgent (cao2025phishagent) use a balanced test set (or 50% base rate). Others such as (liu2023knowledge) and (liu2024less) also use similarly unrealistic base rates.
4.1. Test Dataset
Starting from the clean dataset, we first perform temporal partitioning, followed by leakage filtering to remove test data points that show high similarity to any point in the train set. The pseudocode to generate the test set is shown in Algorithm 2 in Appendix B.
@@ 4.1. Test Dataset
Leakage filter While temporal separation provides some protection against data leakage from the train to the test set, very similar points could still exist between train and test datasets. To mitigate this problem, we traverse the test set and discard points whose TF-IDF cosine distance is less than a specified threshold from any train points. LSH is used to make searching the train set more scalable. Figure 4 in Appendix B shows the temporal distribution of the train and test data sets based on date of collection.
4.2. PhreshPhish Benchmarks
@@
To encourage the development of more robust and generalizable phishing detection models, we provide a suite of benchmark datasets. The benchmark datasets are derived from the test dataset but are further processed to make them more challenging and representative of real-world scenarios. The steps to create the benchmarks are shown in Algorithm 3 in Appendix B.
Diversity enhancement We prune as well as augment the test dataset to make it more diverse and representative. We augment the benign samples with data points corresponding to the top targeted brands to ensure representation of brands not captured by our benign feed. The distribution of targeted brands is estimated across the entire dataset, from which the $k$ most frequent brands are extracted. We perform automated web searches to retreive URLs corresponding to these brands, which are subsequently scraped and integrated into the test dataset. For our dataset, we selected $k=173$, resulting in approximately 1,000 new data points after cleanup.
@@
As part of diversity filter, we prune data points in the test set that have similar URLs or HTMLs. This is applied to both phishing and benign points. This is implemented similar to the leakage filter but points within the test set are compared; any point with neighbors closer than a threshold is dropped.
@@
Difficulty filter Applying the difficulty filter prunes out “easy” points, making the dataset more challenging to classify. There are several approaches to discovering easy points including exploring various data features, determining the most confounding feature values for the output classes, and selecting data points with those values. We take a simpler approach and assume the classification probability of a correctly classified data point can be taken as a proxy of its difficulty. We run the test set through a classifier trained on the training set. Correctly classified phishing points with score greater than $1-\delta$ are considered easy and candidates for pruning. We randomly sample and drop $k\%$ of the entire dataset. Specifically, we set $\delta=0.15$, and $k=10$.
@@
Base rate adjustment The base rate of phishing points in our test set is 46%. Other publicly available test datasets have similar high base rates. However, in practice the base rates a detector observes are much lower, and vary depending on a multitude of factors including data provenance and the model pipeline used in deployment. To represent a wide range of scenarios, we create a large number of benchmark datasets with varying base rates. In particular, we create benchmark datasets with five different base rates varying from 0.05% to 5%. In addition, we also provide the entire benchmark data set, which has a base rate of 24%. To account for variance, multiple instances, $m$, are sampled for each base rate, resulting in 975 total datasets. For lower base rates, the number of phishing points in a benchmark set are miniscule, and hence the variance is high.
@@
Determining $m$: How many instances, $m$, are required, given a base rate, $b$, corresponding to $k$ phishing points in a dataset of size $N$, for the model error estimate, $M_{e}$, to be accurate within a margin of error $e$ (e.g., $\pm 1\%$)? Since model error is an average, we can assume $M_{e}\sim\mathcal{N}(\mu_{M_{e}},\,\sigma_{M_{e}})$ (per central limit theorem). $m$ is sufficiently large when $m\geq\frac{z^{2}\cdot\hat{\sigma_{M_{e}}}^{2}}{e^{2}}$ where $\hat{\sigma_{M_{e}}}^{2}$ is the empirical variance of the model error and $z$ is the standard normal z-score corresponding to a desired confidence level. For each $b$, we sample multiple small sets of instances to compute average empirical variance, which is then used to estimate $m$. This analysis assumes sampling with replacement; sampling without replacement generally requires fewer samples, however, the difference is negligible when $k\ll N$, as is true in our case. Another advantage of sampling without replacement is that all available points are covered. In our experiments, we adopt sampling without replacement with $m$’s higher than the estimated lower bound with 1% margin of error and 95% confidence level. The base rates and the corresponding $m$’s in the benchmark datasets are: 0.05% (586), 0.1% (296), 0.5% (59), 1% (29), 5% (5).
@@
As shown in Figure 3, model performance is highly sensitive to the base rate, with performance improving as the base rate increases. Importantly, most of the strong performance results reported in the literature on this task can be largely attributed to the use of unrealistically high base rates. While Figure 3 empirically shows the dependence of model precision on the base rate, this relationship is supported by theory as well (see Section B.1 in the Appendix).
(a) Linear model
(b) FFN model
(c) GTE model
(d) LLM
Figure 3. Precision-recall curves on the benchmark datasets
5. Baseline Models
@@
Phishing detection is a binary classification problem where the input space $\mathcal{X}$ is the Cartesian product $\mathcal{X}=\mathcal{U}\times\mathcal{H}$ where $\mathcal{U}$ denotes the space of URLs and $\mathcal{H}$ denotes the space of HTML content  [^3] Other input modalities such as host-based features (i.e. domain age, SSL certificate validity, etc.) could be used as well but we exclude them here since their inclusion tends to introduce latency that would be unacceptable for real-time detection.. The goal is to learn a function $f:\mathcal{U}\times\mathcal{H}\rightarrow[0,1]$ that estimates the conditional probability $P(Y=1|x),$ where $Y\in\{0,1\}$ is the binary class label (with $Y=1$ indicating a phishing page and $Y=0$ indicating a benign page) and $x=(u,h)\in\mathcal{X}$. The function is learned from a labeled dataset $\mathcal{D}=\{(x_{i},y_{i})\}^{N}_{i=1}$.
Model 0.05 0.10 0.50 1.0 5.0 24.0 (full)
Linear AP 0.2230 $\pm$ 0.0043 0.3191 $\pm$ 0.0047 0.5825 $\pm$ 0.0050 0.6841 $\pm$ 0.0045 0.8610 $\pm$ 0.0047 0.9551 $\pm$ 0.0028
P@R = 0.90 0.0149 $\pm$ 0.0007 0.0298 $\pm$ 0.0014 0.1223 $\pm$ 0.0062 0.2191 $\pm$ 0.0102 0.5950 $\pm$ 0.0213 0.8957 $\pm$ 0.0075
FFN AP 0.2511 $\pm$ 0.0048 0.3368 $\pm$ 0.0050 0.5798 $\pm$ 0.0058 0.6813 $\pm$ 0.0047 0.8640 $\pm$ 0.0027 0.9607 $\pm$ 0.0018
P@R = 0.90 0.0163 $\pm$ 0.0007 0.0316 $\pm$ 0.0012 0.1319 $\pm$ 0.0044 0.2339 $\pm$ 0.0068 0.6137 $\pm$ 0.0111 0.9045 $\pm$ 0.0038
GTE AP 0.6050 $\pm$ 0.0051 0.6879 $\pm$ 0.0044 0.8354 $\pm$ 0.0035 0.8762 $\pm$ 0.0032 0.9382 $\pm$ 0.0016 0.9779 $\pm$ 0.0020
P@R = 0.90 0.0891 $\pm$ 0.0068 0.1563 $\pm$ 0.0123 0.3736 $\pm$ 0.0289 0.5426 $\pm$ 0.0305 0.8576 $\pm$ 0.0133 0.9746 $\pm$ 0.0037
LLM AP 0.1003 $\pm$ 0.0012 0.1780 $\pm$ 0.0020 0.4839 $\pm$ 0.0038 0.6252 $\pm$ 0.0044 0.8383 $\pm$ 0.0030 0.9216 $\pm$ 0.0034
P@R = 0.90 0.0390 $\pm$ 0.0017 0.0722 $\pm$ 0.0035 0.2606 $\pm$ 0.0141 0.4147 $\pm$ 0.0214 0.7861 $\pm$ 0.0145 0.9352 $\pm$ 0.0235
@@
Table 4. Mean average precision (AP) and mean precision at recall = 0.90 for each benchmark base rate. Values are reported as mean $\pm$ error, where the error reflects the maximum absolute deviation from the 95% bootstrap confidence interval.
We implemented several representative approaches for real-time phishing detecting, including: (1) a linear model trained on n-gram and hand-crafted features; (2) a shallow feedforward neural network (FFN) using the same feature set; (3) a BERT-based model with a classifier head fine-tuned on raw HTML and URLs; and, (4) a large language model (LLM) used in a zero-shot prediction setting. For the linear and FFN models, input features include tokenized character-level n-grams from both the URL and HTML content, as well as hand-engineered features such as the length of the URL, number of nodes in the HTML DOM tree, and other domain-specific features. In contrast, the BERT and LLM-based approaches operate directly on raw textual inputs to learn rich representations. Our objective is not to optimize for state-of-the-art performance, but rather to establish baseline results on the benchmark datasets using common solution approaches. Model architecture and training details for each approach are described in Appendix C.
Linear model The linear model is a simple linear classifier $f(x)=W\cdot x+b$ implemented using LinearSVC from the scikit-learn library (scikit-learn) where parameters $W$ and $b$ are learned by minimizing the L2-regularized squared hinge loss: $\mathcal{L}(W,b)=\frac{1}{N}\sum_{i=1}^{N}(\max(0,1-y_{i}(W\cdot x_{i}+b)))^{2}+\frac{\lambda}{2}\|W\|^{2}_{2}$.
FFN model The FFN model uses the same input representation as the linear model but replaces the linear scoring function with a shallow neural network comprising of a single hidden layer with 32 neurons and ReLU activation. Input features are binarized and subject to frequency-based feature selection to reduce dimensionality. The model outputs a scalar probability estimate $f(x)=\sigma(W_{2}\cdot\texttt{ReLU}(W_{1}\cdot x+b_{1})+b_{2})$. This architecture introduces a small amount of non-linearity while maintaining tractable training and inference costs. Training is performed using binary cross-entropy loss with the Adam optimizer at a learning rate of 0.001.
BERT-based feature learning In this approach, we tokenize the raw, concatenated URL $u\in\mathcal{U}$ and HTML $h\in\mathcal{H}$ and input to a BERT-based encoder model, which generates a hidden contextualized representation, $\mathbf{h}=\texttt{Encoder}(\texttt{tokenize}(u\oplus h))\in\mathbb{R}^{n\times d}$, where $n$ is the number of tokens and $d$ is encoder model dimension. In particular, we use the gte-large pre-trained embedding model (li2023towards), which computes the dense embedding as $x=\frac{1}{n}\sum_{i=1}^{n}h_{i}\in\mathbb{R}^{d}$. We attach a linear classification head for phishing detection, $f(x)=\sigma(W\cdot x+b)\in[0,1]$. We fully fine-tune the model using a binary cross-entropy loss for two epochs on the training dataset.
LLM-based prediction We prompt an LLM (Google’s Gemini-2.5-Pro) with a webpage URL and its procressed HTML to produce a rating indicating the likelihood that the page is a phishing site. The prompt is zero-shot, and the HTML is pruned to remove non-informative elements such as JavaScript, CSS and comments, inspired by (koide2024chatphishdetector). This reduces input tokens and filters out content that is likely to be irrelevant to phishing detection. Further details, including the full prompt, are provided in Appendix C.
5.1. Results and Discussion
Performance Metrics and Evaluation Goals Users downstream of phishing detection systems are typically highly sensitive to false positives – instances where access to legitimate websites is incorrectly blocked. This is quantified by precision, the fraction of times users are correctly blocked. Consequently, commercial products aspire to high values of precision while also maintaining a good recall. Accordingly, we evaluate our models on two performance metrics – the average precision (AP), which approximates the area under the precision-recall curve and the precision at a fixed recall of 0.9 ($P@R=0.9$).
Impact of Base Rate on Performance We evaluate the models on the test dataset (base rate; 46%), the full benchmark dataset (base rate: 24%) and the benchmark datasets with the five base rates ranging from 0.05% to 5%. Table 4 summarizes the results of the four models across the different base rates, while Figure 3 depicts the precision-recall curves. For each rate, the solid line shows the average precision-recall curve of all instances of benchmarks at that base rate. As expected, all models perform strongly on the full benchmark, with most models achieving an AP of more than 95%; the models achieve an even higher AP of greater then 98% on the test set. (For detailed results on the test set, see Section C.3 in the Appendix). The results clearly demonstrate the dependence of model precision on the base rate. As shown in Figure 3, the performance of all models degrades as the base rate decreases. This behavior is theoretically expected (see Section B.1); as the ratio of phish to benign decreases, maintaining high precision at the same recall requires a significantly lower false positive rate (FPR). Further, the degradation is non-uniform across differnt models.
Model Comparison Overall, the GTE model performs the best. In the most challenging scenario (base rate = 0.05), the GTE model achieves a significantly higher AP than the other models. We believe LLM performs poorly compared to the Bert-based embedding model (GTE), primarily since the LLM operates in a zero-shot setting while the embedding model is fully fine-tuned on the training data set. More evaluation results, including precision-recall curves and confusion matrices at multiple threshold values, are provided in Appendix C.
6. Conclusions